### v1

# Large Language Models for Recommendation with Deliberative User Preference Alignment

Abstract.
While recent advancements in aligning Large Language Models (LLMs) with recommendation tasks have shown great potential and promising performance overall, these aligned recommendation LLMs still face challenges in complex scenarios. This is primarily due to the current alignment approach focusing on optimizing LLMs to generate user feedback directly, without incorporating deliberation. To overcome this limitation and develop more reliable LLMs for recommendations, we propose a new Deliberative Recommendation task, which incorporates explicit reasoning about user preferences as an additional alignment goal. We then introduce the Deliberative User Preference Alignment framework, designed to enhance reasoning capabilities by utilizing verbalized user feedback in a step-wise manner to tackle this task. The framework employs collaborative step-wise experts and tailored training strategies for each expert. Experimental results across three real-world datasets demonstrate the rationality of the deliberative task formulation and the superior performance of the proposed framework in improving both prediction accuracy and reasoning quality.
1. INTRODUCTION
Large Language Models are increasingly utilized and aligned with recommendation task through supervised fine-tuning (Bao et al., 2023b; Zhang et al., 2024; Bao et al., 2023a) or direct preference optimization (Chen et al., 2024) for predicting the user feedback. Although these recommendation LLMs (RecLLMs) generally achieve performance gains in different scenarios (Wu et al., 2024; Wang et al., 2024c, b), they can still fail in some complex cases, resulting in valueless even disappointing recommendations. For instance, a RecLLM would present a video on vanilla attention mechanisms to a user who has watched several videos on self-attention and its enhanced variants.
Recent insights from OpenAI regarding the limitations of general LLMs (Guan et al., 2025) suggest that failures of current RecLLMs come from their alignment objective, i.e., optimizing the LLM to directly generate user feedback given the user history and candidate item. This pattern-based learning makes the model susceptible to data biases (Bao et al., 2024) and spurious correlations (Zhang et al., 2023b; He et al., 2023), hindering their ability to generalize to low-frequency or cold-start items. Moreover, this label-oriented optimization objective pushes the LLM to make prediction instantly, discouraging thorough reasoning about user preferences. Consequently, it is crucial to investigate alternative alignment objectives for RecLLMs.
Inspired by the recent advancements in enhancing the reliability of general LLMs through explicit reasoning (Cheng et al., 2025; Ahn et al., 2024; Imani et al., 2023; Li et al., 2024; Xu et al., 2024), we introduce a new Deliberative Recommendation task. This task is designed to train LLMs to engage in reasoning about user preferences and item features prior to predicting user feedback. As illustrated in Figure 1, the RecLLM will explicitly generate a reasoning process, using this process to augment its predictions. To achieve this goal while adhering to commercial restrictions, the key of this formulation is seeking user generated and freely available rationale behind the user feedback as guidance for training the LLM in reasoning generation. Consequently, we incorporate verbalized user feedback such as reviews or conversations, into our task formulation as supervisory signals for reasoning, setting our work apart from previous studies (Tsai et al., 2024; Kim et al., 2024; Long et al., 2024).
To effectively address the Deliberative Recommendation task, it is crucial to decompose the highly personalized and diverse reasoning process (Wang et al., 2022) into distinct steps, and equipping the LLM with specialized reasoning capabilities for each step. In this way, we highlight three fundamental capabilities: 1) Preference Distillation, which analyzes the user history to identify aspect-level user preferences (Guan et al., 2019) and examines existing verbalized user feedback on the candidate item to recognize positive and negative features; 2) Preference Matching, which matches distilled user preferences with item features and generates rationales for why the user might like or dislike the candidate item; and 3) Feedback Prediction, which predicts the user feedback with consideration of the generated raionales.
In this light, we propose a Deliberative User Preference Alignment (DeliRec) framework to teach LLMs with the recognized reasoning capabilities with the guidance of verbalized user feedback (e.g., user reviews (Hasan et al., 2024)). As illustrated in Figure 2, DeliRec employs three collaborative experts, each dedicated to a specific reasoning step, to hone step-specific reasoning capabilities. To reduce the computational and memory cost, these experts are implemented using separate QLoRA (Dettmers et al., 2023) adapters on the same LLM. Recognizing that each reasoning step necessitates extracting distinct signals from the verbalized user feedback, each expert is equipped with a uniquely tailored training strategy. Taking rating prediction (Liu et al., 2019a) as an example, we conduct extensive experiments across three real-world datasets, validating the effectiveness of DeliRec regarding prediction accuracy and reasoning quality, and the rationality of the framework design. The code and datasets are available at https://anonymous.4open.science/r/OurDeliRec-F4B5.
To summarize, our contributions are threefold:
We formulate the Deliberative Recommendation task, which pursues LLMs conducting reasoning before making prediction by learning from verbalized user feedback.
We propose a Deliberative User Preference Alignment framework, which achieves the reasoning process with step-wise experts associated with specifically designed training strategies.
We conduct extensive experiments on three datasets, validating the effectiveness and rationality of the proposed DeliRec framework, showing the potential of slow thinking in recommendation.
2. TASK FORMULATION
Deliberative Recommendation. We formulate the task of deliberative recommendation, which performs thoughtful reasoning regarding user preference and item features before predicting user feedback as illustrated in Figure 1. To guide the reasoning process, we propose incorporating verbalized user feedback for LLM optimization. In this work, we adopt user reviews due to fine-grained user preferences described in reviews (Hasan et al., 2024). Besides, we investigate deliberative recommendation on the classic rating prediction task, following prior work (Tsai et al., 2024; Kim et al., 2024).
Formally, let denote a set of users and a set of items. The interaction between user and item can be represented as , where is the rating of user for item , and is the corresponding review, which exists in many datasets (Harper and Konstan, 2016; Luo et al., 2022). For a user-item pair , the historical interactions of user and item are defined as:
where represents items previously rated by user , and denotes users who rated item . Given and , the LLM is tasked to predict the rating after performing explicit reasoning:
where denotes the reasoning process and is the predicted rating.
Considering the reasoning process is highly personalized and diverse across users, we decompose the complex reasoning into multiple steps, as shown in Figure 2, to achieve different objectives: Preference Distillation, Preference Matching, and Feedback Prediction. To emphasize the alignment between the reasoning process and the user’s true preference, we incorporate the target item review of user for item as additional supervision during training. Notably, it is excluded in testing to prevent data leakage.
Justification for explicit reasoning in RecLLMs. Incorporating explicit reasoning into RecLLMs holds significant potential, despite it potentially increases the prediction time and computational cost. Specifically, it enables RecLLMs to: 1) fully leverage the reasoning capabilities of LLMs to improve prediction accuracy and reliability; 2) enhance interpretability through step-by-step reasoning as explanations, thereby enhancing user trust; 3) possibly facilitate user-controllable recommendation for human-AI collaboration (Yang et al., 2024), since users can adjust the reasoning process to generate new recommendations. Notably, explicit reasoning is particularly beneficial in scenarios where real-time recommendation requirements are less critical, but accuracy is paramount, such as E-commerce product, movie, and medication recommendations (Ali et al., 2023). These items typically involve long usage periods and require careful deliberation before purchase or viewing. Furthermore, as LLMs efficiency continually improves (Lin et al., 2024), the computational costs will decrease accordingly, enhancing the feasibility of wide applications.
3. METHOD
In this section, we present DeliRec with three experts to collaboratively perform multi-step reasoning before recommendation. To supervise the reasoning process, DeliRec incorporates user reviews, directly aligning the reasoning with user preferences rather than only inferring preferences through users’ numerical or binary feedback (e.g., ratings and clicks). We then detail the three experts, i.e., summarizer, reasoner, and predictor, with their fine-tuning strategies. An overview of DeliRec is illustrated in Figure 2.
3.1. Summarizer for Preference Distillation
DeliRec leverages a summarizer to distill the user’s preferences and the item’s positive and negative features from their historical reviews. It filters out noise from the reviews, captures nuanced user preferences, and enriches item features with the LLM’s inherent knowledge. The distillation output is formatted as an aspect-preference summary, i.e., concise and representative keywords that encapsulate essential user preferences and item features, providing a reasoning foundation for the subsequent reasoning process.
Summarizer Instantiation. Formally, given a historical interaction between user and item , the summarizer will generate an aspect-preference summary. We instantiate the summarizer by fine-tuning a QLoRA adapter (Dettmers et al., 2023) on an LLM and the summarizer’s prompt is illustrated in the following, which uses the Music dataset as an example.
Summarizer Prompt Task: Summarize the reasons behind the given rating of a Music based on the customer review. Music: Rating: Review: Analyze the above customer review for the Music and summarize the reasons behind the given rating of . Please consider the positive and negative aspects mentioned in the review and provide the keywords of reasons and user preference elements.
The user’s preferences and the item’s features are outputted as an aspect-preference summary. A template example is as follows:
Positive Aspects: Catchy Melody, Unique Instrumentation, … Negative Aspects: Repetitive Lyrics, Overuse of Autotune, … User Preference Elements: Harmony, Emotional Resonance, …
Summarizer Fine-tuning. Considering ChatGPT (OpenAI, 2023) has shown strong performance in various summarization tasks (Basyal and Sanghvi, 2023), we employ closed-source ChatGPT as a teacher model to generate aspect-preference summaries, guiding the learning of open-source LLMs. With these summaries, we optimize the QLoRA adapter of the summarizer by SFT (Ouyang et al., 2022).
Offline Summarization. Given a summarizer, we can store the generated aspect-preference summaries offline for all the historical interactions between users and items. This step eliminates the need for repeated preference distillation, thereby reducing computational overhead and enhancing efficiency.
3.2. Reasoner for Preference Matching
Given all the aspect-preference summaries of a user and an item, DeliRec adopts a reasoner to evaluate their matching degree, generating reasons for why the user might like or dislike the item.
Reasoner Instantiation. Formally, given a user-item pair , we aggregate the aspect-preference summaries from their historical interactions and in chronological order into a single prompt, tasking the reasoner to measure the preference matching between and . The reasoner is also implemented by another QLoRA adapter on the same LLM. An example of the reasoning prompt is shown below:
Reasoner Prompt ### User Review History ### organized as below 1. Title of Item 1 Positive Aspects: [Aspect 1], [Aspect 2], … Negative Aspects: [Aspect 1], [Aspect 2], … User Preference Elements: [Preference 1], [Preference 2], … 2. Title of Item 2 Positive Aspects: [Aspect 1], [Aspect 2], … Negative Aspects: [Aspect 1], [Aspect 2], … User Preference Elements: [Preference 1], [Preference 2], … … ### Item Review History by Other Users ### organized in the same format as above Task: Analyze whether the user will like the new Music based on the user’s preferences and the item’s features. Provide your rationale in one concise paragraph.
Reasoner Fine-tuning. Given a historical interaction between user and item , we propose using target item review to guide the reasoner learning, because users would explicitly explain their preferences in verbalized reviews. However, directly using reviews for supervision suffers from noise issues, since user reviews often contain low-quality comments or preference-irrelevant details (see evidence in Section 4.3.2). Thus, we employ a generation-then-filter strategy: using ChatGPT first generates diverse reasons why user might like or dislike item , and a reward model then selects the high-quality reasons based on target item review . The selected high-quality reasons are used for reasoner learning through SFT.
Reason Generation. As outlined in Algorithm 1, given a user-item pair, we repeatedly sample candidate reasons from ChatGPT based on the reasoner prompt and evaluate reasons using the reward model until either a high-quality reason is obtained or the maximum iteration count is reached. Particularly, to avoid repeatedly generating reasons that contradict the user’s binary preference, we incorporate the user’s actual rating as a hint, except for the initial generations. This is implemented by appending the following hint statement to the reasoner prompt:
Hint: The user actually rated the item stars. The star ranges from 1 to 5, with 5 being the best. Use the hint but don’t mention the user’s rating in your response.
Reason Filter. We introduce a reward model to filter candidate reasons based on their alignment with the target item review . Considering the alignment, the reward model assigns a score of 1 to high-quality reasons and 0 to low-quality reasons. Using this reward model, we can iteratively generate high-quality reasons using ChatGPT for the reasoner fine-tuning. The detailed training and inference of the reward model can be found in Section 3.4.
3.3. Predictor for Feedback Prediction
DeliRec finally introduces a predictor for user feedback prediction. In this work, we investigate the classic rating prediction task (Zheng et al., 2017). The input for the predictor is the reasoning process generated by the summarizer and reasoner. In addition, considering that the user’s historical ratings serve as a valuable reference for rating prediction (Koren et al., 2009), we also include ratings as input. The user’s actual rating for the target item is used for supervision.
Predictor Instantiation. Given a user-item pair with historical interactions , we first aggregate the aspect-preference summaries from and in the same format as in the reasoner prompt, supplemented with historical ratings (see the following predictor prompt). Next, following previous work (Kim et al., 2024), we include the average ratings from all historical interactions in the prompt. Finally, the reason generated by the reasoner is appended to the Reason Placeholder. The final prompt is shown below to predict the user ratings. To implement the predictor, we apply another QLoRA on the LLM and utilize the rating for SFT.
Predictor Prompt ### User Review History ### is organized in the same format as in the reasoner prompt, except that historical ratings are additionally included. 1. Title of Item 1, Rating: 5.0 Positive Aspects: [Aspect 1], [Aspect 2], … Negative Aspects: [Aspect 1], [Aspect 2], … User Preference Elements: [Preference 1], [Preference 2], … … ### Item Review History by Other Users ### organized in the same format as above ### Average Past Ratings ### User’s Average Rating (all previous ratings): 4.5 Item’s Average Rating (all ratings by other users): 3.7 ### Personalized Recommendation Analysis ### This section is referred to as Reason Placeholder, where we place the generated reason here. Task: Based on the above information, please predict the user’s rating for , (1 being the lowest and 5 being highest, directly give the rating without other content.) [Output Format] Predicted Rating:[Rating between 1 and 5]
Logit-weighted Decoding. We adopt logit-weighted decoding for the final prediction, which transforms the predicted rating scores from discrete integer values to continuous scores—a method proven to offer greater precision and adopted in previous work (Kim et al., 2024).
Specifically, when decoding the rating token, the LLM first obtains the logit values for each possible rating token (e.g., the logits for tokens “1”, “2”, “3”, “4”, and “5”), denoted as (assuming the rating range is from one to five). These logits are then normalized into probabilities for each rating using the softmax function: . The final prediction is calculated as the expected value: , where each possible rating is weighted by its corresponding probability.
Efficient Rating Prediction. Since the LLM is instructed to output ratings in a fixed format, starting with the phrase “Predicted Rating:” as specified in the predictor prompt, we can skip decoding these predictable and invariant words. We directly decode the final rating using logit-weighted decoding. Only decoding the single rating token significantly reduces inference time and costs.
3.4. Reward Model for Reason Judgment
We use a reward model to select high-quality reasons for the reasoner fine-tuning. Given candidate reasons and the corresponding target item reviews, the reward model assigns a score of 1 if the reason is high-quality and aligns with the target item review; otherwise, it assigns 0.
Reward Model Training. We fine-tune another QLoRA on the LLM as the reward model, which is trained to predict the user ratings based on the user review. The input is the predictor prompt template, where the Reason Placeholder is filled by the target item review rather than the generated reasons. The reward model is trained to extract the user’s true preferences from the target item review to enhance prediction accuracy.
Separate Training Data. To prevent the reward model from predicting ratings solely by memorizing training data rather than capturing the user’s true preferences from target item reviews, we separate the training data for the reward model and the reasoner. This separation ensures that the reward model is evaluated on unseen data, enabling a more reliable assessment of reason quality.
Reason Judgment. Since the reward model is trained to capture the user’s true preferences from the target item’s review, a reason that better aligns with the user’s review can achieve superior prediction accuracy, as evidenced in Section 4.3.3. Based on this insight, we define an evaluation score as:
where represents the ground-truth rating for user and item . and denote the predicted ratings by the reward model based on the target item review and a generated reason. And measures the prediction difference:
If , it suggests that the reason surpasses the raw review in rating prediction performance.
If , a smaller indicates a closer performance between the reason and review, suggesting superior alignment.
By introducing a threshold , we identify all reasons of as high-quality ones with the evaluation score as 1. Treating as a hyperparameter can balance the trade-off between the quantity and quality of the reasons.
4. EXPERIMENT
In this section, we conduct a series of experiments to answer the following research questions:
RQ1: How does DeliRec perform compared to other baseline methods in terms of accuracy in predicting user feedback (i.e., ratings) and the quality of generated reasons?
RQ2: Why is the multi-step reasoning strategy in DeliRec essential, and how does it compare in effectiveness to one-step reasoning strategies?
RQ3: How effective is our approach (the reason generation method and proposed reward model) in incorporating user-verbalized preference feedback?
4.1. Experimental Settings
4.1.1. Datasets.
We conduct experiments on three widely-used real-world datasets:
Amazon Music (Music): This refers to the “Digital Music” subset of the well-known Amazon Product dataset111https://cseweb.ucsd.edu/~jmcauley/datasets/amazon/links.html., which records rich user reviews, ratings, and textual information about items, such as titles, across a broad range of product categories, on the Amazon platform. We refer to this dataset as “Music” for short.
Amazon Book (Book): This refers to the “Book” subset of the Amazon Product dataset, shortened “Book”.
Yelp: This refers to the Yelp Open dataset222https://business.yelp.com/data/resources/open-dataset/., which includes user reviews, ratings for businesses such as restaurants and retail shops, as well as textual information about the businesses. It is widely used in recommendation tasks (Qiu et al., 2021).
We use the entire Music dataset for experiments, while for the Book and Yelp datasets, we utilize only a subset due to their large size. For the Book dataset, we use data from the last two months, and for the Yelp dataset, we use data from the last six months. For each dataset, we split it into training, validation, and test sets based on the timestamps of interactions, ensuring that test interactions occur after all training and validation interactions to prevent information leakage (Ji et al., 2023). Regarding data filtering, following prior work (Liu et al., 2019b), we adopt a 5-core setting to filter the data and exclude cold-start users and items—those not appearing in the training set—from the validation and test sets. The statistical details of the processed dataset are provided in Table 1.
| Dataset | #Train | #Valid | #Test | #User | #Item |
| Music | 43,071 | 3,271 | 1,296 | 4,183 | 2,660 |
| Book | 71,972 | 6,144 | 5,541 | 13,863 | 13,515 |
| Yelp | 51,497 | 4,757 | 4,328 | 8,453 | 13,426 |
4.1.2. Evaluation Metrics.
To evaluate the accuracy of rating prediction, we use two standard metrics (Kim et al., 2024; Liu et al., 2023): Mean Absolute Error (MAE) and Root Mean Square Error (RMSE).
To evaluate reasoning quality, we employ BLEURT (Sellam et al., 2020) and GPTScore (Wang et al., 2023a) to measure the semantic alignment between the generated reasoning and the target item review. BLEURT333https://github.com/google-research/bleurt. is an evaluation metric that leverages contextual embeddings to assess the semantic similarity between text pairs, providing a fine-grained measure of alignment. GPTScore, computed based on GPT-4o444https://chatgpt.com/?model=gpt-4o., analyzes the semantic alignment of the given text pairs and assigns a score between 0 and 100, with higher scores indicating better alignment. These metrics are commonly adopted in explainable recommendation tasks (Ma et al., 2024).
4.1.3. Baselines.
To comprehensively evaluate performance, we compare DeliRec against three categories of recommendation methods: traditional methods (MF), review-based recommendation methods (DeepCoNN, NARRE, DAML), and LLM-based methods (GPT-4o, Rec-SAVER, EXP3TR).
MF (Koren et al., 2009): Matrix Factorization is a classical collaborative filtering method that predicts user ratings based solely on the historical rating matrix.
DeepCoNN (Zheng et al., 2017): This method employs a Convolutional Neural Network (CNN) model to jointly learn item properties and user behaviors from review text to assist in rating prediction.
NARRE (Chen et al., 2018): This method employs an attention mechanism to prioritize the most useful reviews, enabling better extraction of user and item features to enhance recommendations.
DAML (Liu et al., 2019a): This method models the interaction between user and item review documents to improve the representations of both users and items.
Rec-SAVER (Tsai et al., 2024): This method supplies LLMs with the target item’s metadata (e.g., title, categories, description) and the user’s historical interactions (e.g., metadata of previously interacted items, user ratings, and raw reviews), and then instructs the LLMs to generate their reasoning process before providing the prediction.
EXP3RT (Kim et al., 2024): This method first constructs user and item profiles from historical reviews and ratings using LLMs. These profiles are then leveraged to reason about user preferences and predict potential ratings in a single step.
4.1.4. Implementation Details.
For review-based methods, we adopt the implementation555https://github.com/ShomyLiu/Neu-Review-Rec. from previous work (Liu et al., 2019b). For all reasoning-enhanced LLMRec methods (i.e., Rec-SAVER, EXP3RT and our DeliRec), we use GPT-3.5-turbo666https://platform.openai.com/docs/models#gpt-3-5-turbo. as the teacher model and fine-tune the LLama3-8B model777https://www.llama.com/docs/model-cards-and-prompt-formats/meta-llama-3/. using the QLoRA technique. To accelerate the training and inference process, we leverage the Unsloth888https://github.com/unslothai/unsloth. framework.
To reduce computational costs, we randomly sample 12,000 user-item pairs from the training data to construct the instruct data for LLMs learning, while conventional methods use the full train dataset. This sample size is sufficient to achieve strong performance, as demonstrated in our main results. All reasoning-enhanced LLMRec methods are trained on the same 12,000 user-item pairs. For training the reward model, we use an additional 8,000 user-item pairs, ensuring no overlap with the instruct data. The hyper-parameter reward threshold ( ) is set to 0.1, 0.2, and 0.04 for Music, Book, and Yelp, respectively. More details can be found in our code.
4.2. Main Results (RQ1)
We compare our DeliRec method with baseline methods on the accuracy of user feedback (rating) prediction (Predictor output) and the quality of generated reasons (Reasoner output).
Accuracy of User Feedback Prediction. The comparison of user feedback prediction performance is summarized in Table 2. As shown in the table, for the MAE metric, our method outperforms all datasets, achieving an average MAE reduction of approximately 0.02 compared to the best baseline results. For the RMSE metric, our method achieves the best performance on the Amazon Music and Amazon Book datasets, while delivering comparable results to the best baseline on the Yelp dataset. These superior performances of DeliRec confirm that incorporating reasoning about user preference and item features before making the final prediction, i.e., applying deliberation, is effective in enhancing RecLLM performance, thereby validating the rationality of our formulation of deliberative recommendation. Notably, although Rec-SAVER and EXP3RT also incorporate reasoning, their reasoning is performed in a single step, unlike our approach, which involves multiple steps—preference distillation, matching, and prediction. This lack of true “deliberation” in a single step may lead to their relatively poorer performance.
Quality of Generated Reasons. We next analyze the quality of the reasons generated by the Reasoner in DeliRec, compared to two reasoning-aware baselines (i.e., Rec-SAVER and EXP3RT). Table 3 summarizes the results. As shown by the results of the two metrics, whether compared with target item reviews using the BLEURT metric or evaluated using the more powerful GPT model (corresponding to GPTScore metric), the reasons generated by our method significantly outperform those of the two reasoning-aware baselines. Our method achieves an average relative improvement of 7.9% on the GPTScore metric and 7.0% on the BLEURT metric across all datasets. The superiority of our method’s reasoning can be attributed to the fact that we teach LLMs with recognized reasoning capabilities, guided by verbalized user feedback (e.g., user reviews (Hasan et al., 2024)), while the baselines fail to leverage this guidance.
| Type | Method | Music | Book | Yelp | |||
| MAE | RMSE | MAE | RMSE | MAE | RMSE | ||
| CF-based | MF | 0.6188 | 0.8142 | 0.6277 | 0.8565 | 0.7980 | 1.0711 |
| Review-based | DeepCoNN | 0.6034 | 0.8057 | 0.6221 | 0.8403 | 0.8312 | 1.0665 |
| NARRE | 0.5799 | 0.7881 | 0.6242 | 0.8435 | 0.8177 | 1.0785 | |
| DAML | 0.5703 | 0.7848 | 0.6214 | 0.8371 | 0.7964 | 1.0405 | |
| LLM-based | GPT-4o | 0.7438 | 1.1069 | 0.7591 | 1.1558 | 0.8766 | 1.3005 |
| Rec-SAVER | 0.6463 | 0.9262 | 0.6645 | 0.9356 | 0.8295 | 1.1282 | |
| EXP3RT | 0.5608 | 0.8385 | 0.6135 | 0.9370 | 0.8306 | 1.2311 | |
| Ours | DeliRec | 0.5442 | 0.7722 | 0.6029 | 0.8345 | 0.7586 | 1.0418 |
| Method | Music | Book | Yelp | |||
| GPT | BLEURT | GPT | BLEURT | GPT | BLEURT | |
| Rec-SAVER | 75.60 | 0.3652 | 72.45 | 0.4233 | 66.43 | 0.4102 |
| EXP3RT | 76.22 | 0.3840 | 73.60 | 0.4373 | 64.28 | 0.4275 |
| DeliRec | 80.53 | 0.4067 | 77.31 | 0.4731 | 72.70 | 0.4565 |
4.3. In-depth Analysis
In the following sections, we present an analysis of the impact of multi-step reasoning and verbalized preference feedback in DeliRec, followed by the evaluation of DeliRec’s inference cost. Lastly, we provide detailed cases comparing the reasoning generated by DeliRec with those from baseline methods to demonstrate its superior alignment with the user’s true preferences.
4.3.1. Impact of Multi-step Reasoning (RQ2)
To evaluate the impact of multi-step reasoning in DeliRec, we first conduct an ablation study to analyze the necessity of the component in each step (i.e., Summarizer, Reasoner and Predictor) and then assess the performance changes by replacing the multi-step reasoning mechanism with two alternative single-step reasoning strategies.
Necessity of Each Step. An ablation study on DeliRec is conducted by removing each step to assess its contribution. Specifically, we make the following changes in each experiment: 1) w/o Step 1 (i.e., Summarizer): Directly using user historical reviews to replace the aspect-preference summaries as input for Step 2 and Step 3; 2) w/o Step 2 (i.e., Reasoner): Remove the Reason Placeholder section from the Step 3 Prompt and instruct the LLMs to directly predict the rating based solely on the aspect-preference summaries and historical ratings; 3) w/o Step 3 (i.e., Predictor): Combine Step 2 and Step 3 into a single step, asking the LLMs to simultaneously generate the match analysis and predicted ratings based on the aspect-preference summaries and the user’s historical ratings.
We conduct this study on Amazon Music and Book datasets and present the results in Figure 3, from which we make the following observations:
The removal of each step results in an increase of the rating prediction errors on both Music and Book datasets, as reflected by higher MAE and RMSE values in the figure.
The MAE and RMSE increase significantly on both datasets after removing Step 3, underscoring the importance of separately training match analysis and rating prediction. One potential reason is that simultaneously training the LLM to complete the reasoning and rating tasks imposes significant challenges, resulting in degraded performance.
The removal of Step 2 results in a relatively slight increase in RMSE but a significant increase in MAE on both datasets. This discrepancy can be attributed to the fact that RMSE is overly sensitive to extreme errors (e.g., predicting 1 for a true rating of 5). Without match analysis, LLMs can still predict a roughly correct rating based on the user’s aspect-preference summaries and historical ratings, which helps avoid extreme errors and results in only a slight increase in RMSE. In contrast, MAE assigns equal weight to all errors and is better at capturing finer-grained inaccuracies. The significant increase in MAE highlights the critical role of Step 2 in enabling the LLMs to predict user feedback more effectively on a fine-grained scale.
One-step v.s. Multi-step. To verify the effectiveness of the multi-step reasoning strategy in DeliRec, we compare it with two alternative strategies: (1) One-step reasoning, which generates both an entangled reasoning process and the ratings in a single step, and (2) Chain-of-Thought (CoT) reasoning, which generates the outputs of our three reasoning steps in one single prompt. We prepare the datasets separately based on the two above strategies and train the LLMs accordingly. In Figure 4, we summarize the results and make the following observations: 1) Both one-step reasoning and CoT reasoning strategies lead to a notable increase in prediction errors on both Music and Book datasets, as reflected by the higher MAE and RMSE values. 2) In contrast, the performance of the CoT reasoning strategy is inferior to that of the simpler one-step reasoning strategy. This may be due to the complexity of multiple tasks involved in the CoT reasoning strategy, which poses significant challenges for LLMs to effectively process simultaneously.
4.3.2. Effect of Reason Generation Approaches (RQ3)
To validate the effectiveness of our reason generation approach in Step 2 of DeliRec, we compare it against four alternative approaches: 1) Without reason: Remove Step 2 from DeliRec entirely and instruct the LLMs to directly predict the rating based solely on the aspect-preference summaries and historical ratings, as done in the ablation study; 2) Review-guided reasons: Use the target item reviews as direct supervision for the LLMs’ learning process in Step 2; 3) Review-inferred reasons: Use ChatGPT to generate reasons inferred from the target item reviews and employ these generated reasons as supervision for Step 2; 4) History-inferred reason: Use ChatGPT to generate reasons inferred from the aspect-preference summaries using the Step 2 Prompt and directly use them as supervision for Step 2 without applying filtering through our reward model.
In Table 4, we present the results and make the following observations: 1) All four alternative approaches perform worse than our proposed method on both Music and Book datasets, highlighting the advantages of our proposed method in collecting high-quality reasons for training LLMs. 2) The performance of Review-inferred Reason is worse than Without Reason on both datasets in MAE and RMSE. This indicates that directly using reviews to generate reasons can have a negative impact. A possible explanation is that our objective is to train LLMs to perform reasoning based on historical interactions without relying on target item reviews. Using reasons derived directly from these reviews introduces the risk that the reasoning logic may rely on preferences not reflected in the historical data. Such inconsistencies disrupt the training process and ultimately result in suboptimal performance. 3) History-inferred Reason also performs worse than “Without Reason” on most metrics. This indicates that directly supervising with GPT-generated reasons without refinement is ineffective, underscoring the importance of using verbalized preference feedback as supervision.
| Method | Music | Book | ||
| MAE | RMSE | MAE | RMSE | |
| Without Reason | 0.5529 | 0.7742 | 0.6225 | 0.8394 |
| Review-guide Reason | 0.5825 | 0.7812 | 0.6376 | 0.8390 |
| Review-inferred Reason | 0.5723 | 0.7816 | 0.6306 | 0.8420 |
| History-inferred Reason | 0.5640 | 0.7744 | 0.6305 | 0.8376 |
| DeliRec | 0.5442 | 0.7722 | 0.6029 | 0.8345 |
4.3.3. Effectiveness of Reward Model (RQ3)
The reward model should be capable of assessing whether the generated reasons align with the ground-truth preferences (as reflected in the target item reviews) by utilizing these reasons for rating prediction. We conduct experiments to verify whether the reward model successfully acquires this capability. Specifically, we construct different types of reasons with known similarity levels to target item reviews to compare the prediction accuracy. We include the four types of reviews defined in Section 4.3.2, along with an additional type: Hint-Inferred Reason. This type generates reasons using ChatGPT, based on aspect-preference summaries and our reasoner’s prompt, with the user’s actual rating included as a hint. Based on their similarity to ground-truth reviews, the types of reasons are ranked from highest to lowest as follows: Review-Guided, Review-Inferred, Hint-Inferred, and History-Inferred Reason. Table 5 summarizes the reward model’s prediction accuracy using each type of reason. The results show that the prediction accuracy ranking aligns with the known similarity ranking, verifying that the reward model effectively evaluates the quality of the generated reviews.
| Music | Book | |||
| MAE | RMSE | MAE | RMSE | |
| Without Reason | 0.6798 | 0.9359 | 0.6186 | 0.8298 |
| Review-guided Reason | 0.294 | 0.4527 | 0.3329 | 0.4597 |
| Review-inferred Reason | 0.4238 | 0.568 | 0.4214 | 0.5394 |
| Hint-inferred Reason | 0.4879 | 0.6876 | 0.5174 | 0.7137 |
| History-inferred Reason | 0.7257 | 1.071 | 0.6842 | 0.9686 |
4.3.4. Cost Analysis
To evaluate the efficiency of DeliRec, we compared its inference time and generation cost with those of reasoning-aware baselines (Rec-SAVER and EXP3RT). As shown in Table 6, despite dividing the reasoning process into three steps, DeliRec achieves comparable average inference time and generates a similar number of tokens as the baselines. Specifically, it requires approximately 6 seconds and generates an average of 148 tokens per prediction. This efficiency is primarily attributed to two factors: 1) the preference distillation step in DeliRec is preprocessed in advance, as discussed in Section 3.1, leaving only the last two steps to be executed during inference; and 2) DeliRec only decodes a single token during the final step, as explained in Section 3.3. These designs effectively minimize computation costs.
| Method | Avg. Inference Time (s) | Avg. Tokens Generated |
| DeliRec | 5.86 | 147.78 |
| Rec-SAVER | 6.43 | 175.59 |
| EXP3RT | 5.62 | 150.74 |
4.3.5. Case study
We conduct a case study to examine whether the reasons generated by DeliRec better align with true user preferences, as reflected in the target item reviews. We compare the reasoning results of DeliRec with those of EXP3RT and Rec-SAVER. Figure 5 illustrates an example. In this case, the target item review indicates that the user values the book for its practicality and helpfulness. DeliRec accurately identifies these features and aligns them with the user’s preferences, as highlighted in green. In contrast, EXP3RT incorrectly attributes the user’s preference to the storyline and characters. Rec-SAVER also fails to recognize both the item’s key features and the user’s actual preferences, despite predicting that the user would like the book.
5. RELATED WORK
In this work, we focus on Deliberative Recommendation, empowering RecLLMs with explicit reasoning abilities and optimizing LLMs through verbalized user feedback. This is highly related to LLM-based recommendation, deliberative alignment, and review-based recommendation.
5.1. LLM-based Recommendation
Recent research has shown growing interest in RecLLMs (Wang et al., 2024b), with many studies exploring the use of LLMs to predict user feedback through in-context learning (Liu et al., 2023; Dai et al., 2023; Sanner et al., 2023) and fine-tuning (Bao et al., 2023b; Zhang et al., 2024, 2023a). However, these approaches often instruct LLMs to generate predictions directly without disclosing their intermediate reasoning steps, resulting in a limited utilization of the amazing reasoning capabilities of LLMs. To address this limitation, some recent studies explored leveraging LLMs’ reasoning capabilities for recommendation tasks through various prompting strategies, such as chain-of-thought (CoT) (Wei et al., 2022) and self-reflection (Wang et al., 2023b). For instance, DRDT (Wang et al., 2023b) encourages LLMs to perform sequential recommendation in a divergent thinking manner by prompting them to analyze user preferences from multiple aspects, while simultaneously reflecting on their analysis using a critic prompt (Wang et al., 2023b). GOT4Rec (Long et al., 2024) applies the graph of thought strategy to prompt LLMs to reason through three different directions. However, these in-context learning methods are inherently constrained by the models’ existing capabilities, making LLMs often struggle to handle recommendation tasks effectively without task-specific fine-tuning (Bao et al., 2023b).
Some studies have explored fine-tuning LLMs to improve their reasoning ability specifically for RecLLMs (Kim et al., 2024; Tsai et al., 2024; Bismay et al., 2024; Wang et al., 2024d). For example, RecSAVER utilizes reasoning generated by a larger LLM as ground truth to fine-tune a smaller LLM, enabling the latter to learn reasoning patterns tailored for recommendation tasks (Tsai et al., 2024). EXP3RT (Kim et al., 2024) employs a similar fine-tuning strategy but introduces an additional preparatory step: to construct user and item profiles based on their review histories first. Leveraging these constructed profiles, EXP3RT reasons about user preferences and predicts possible ratings in the same single step. However, these fine-tuning methods face two main limitations. First, they compress the complex reasoning process into a single step and train it jointly, which makes it challenging for LLMs to effectively learn. In contrast, DeliRec decomposes the reasoning process into multiple steps, training each step independently for better performance. Second, they lack fine-grained supervision in training data selection, leading models to learn the reasoning patterns misaligned with users’ real preferences. DeliRec addresses this limitation by using verbalized user feedback as supervision, resulting in reasons that more accurately reflect users’ preferences and enhance the multi-step reasoning abilities of RecLLMs.
5.2. Deliberative Alignment
Existing work has demonstrated the superiority of LLMs combined with multi-step reasoning (Wang et al., 2024a), with OpenAI’s O1 showcasing remarkable reasoning capabilities (OpenAI, 2024). Meanwhile, Deliberative Alignment (Guan et al., 2024), a training paradigm that directly teaches LLMs human-written safety specifications and trains them to explicitly reason about these specifications before responding, has garnered widespread attention in the LLM safety domain.
However, there are significant differences between this and the Deliberative Recommendation task: 1) in the safety domain, alignment specifications are relatively well-defined and consistent, whereas the reasoning process behind user feedback is highly complex, with significant personalization and diverse preferences. Achieving deliberative user preference alignment at the individual level is thus a far more challenging task. 2) Given the complexity of the Deliberative Recommendation task, relying solely on numerical user feedback (e.g., ratings) or pairwise preference data for LLM learning, is insufficient. We propose leveraging verbalized user feedback to supervise reasoning for superior alignment, setting it apart from previous approaches to Deliberative Alignment.
5.3. Review-based Recommendation
Review-based recommendation systems have been widely studied recently to address tasks such as rating prediction (Qiu et al., 2021; Zhao et al., 2019; Kim et al., 2016; Yang et al., 2023; Shuai et al., 2022). These methods typically adopt a dual tower architecture that employs a user encoder and an item encoder to capture the semantics of reviews as user and item embeddings. Then, the user and item embeddings are combined through a fusion layer to derive joint representations, which are passed to a rating prediction layer to produce the final output (Hasan et al., 2024). To obtain better embeddings of users and items from reviews, many advanced techniques have been explored. For instance, DeepCoNN (Zheng et al., 2017) employs convolutional neural networks (CNNs) to extract semantic information from historical reviews, achieving excellent results. Furthermore, attention mechanisms (Bahdanau et al., 2015) are also introduced at both the word and review levels to identify critical words and reviews that are most significant (Liu et al., 2019b; Tay et al., 2018; Wu et al., 2019). For example, NARRE (Chen et al., 2018) employs attention mechanisms to prioritize reviews that are most useful for improving embedding learning. While these methods leverage reviews to represent user preferences implicitly with embeddings, resulting in the lack of interoperability. In contrast, DeliRec generates the reasoning process for each step in explicit textual form, offering clearer and more interpretable explanations of user preferences. Additionally, most review-based recommendation methods primarily leverage reviews as input features for generating embeddings. However, DeliRec also utilizes reviews as supervisory signals on the output side, aligning the LLM’s reasoning with the user’s true preferences.
6. CONCLUSION AND FUTURE WORK
In this work, we addressed the limitations of existing RecLLMs in complex scenarios by introducing a novel Deliberative Recommendation task, which emphasizes explicit reasoning before predicting user feedback. To achieve Deliberative Recommendation, we proposed the DeliRec framework, which enables multi-step reasoning via three collaborative experts with three core reasoning capabilities: Preference Distillation, Preference Matching, and Feedback Prediction. We aligned the reasoning process with users’ true preference by using verbalized user feedback, i.e., reviews. Through extensive experiments on three real-world datasets, DeliRec demonstrated superior performance in both prediction accuracy and reasoning quality, highlighting the significance of slow thinking in recommendation tasks.
Despite its promising results, DeliRec is an initial attempt on Deliberative Recommendation, leaving many future directions. First, we only adopt user reviews as verbalized user feedback. Exploring multi-turn and diverse forms of verbalized user feedback, e.g., conversations, may further enhance the reasoning process. Second, although DeliRec achieves comparable or even lower inference costs than Rec-SAVER and EXP3RT, there is still significant room to improve its efficiency. Future work could explore designing LLM acceleration algorithms tailored to Deliberative Recommendation. Third, it is promising to design an interactive learning paradigm between DeliRec and users.

### v2

@@
-Large Language Models for Recommendation with Deliberative User Preference Alignment
+Reason4Rec: Large Language Models for Recommendation with Deliberative User Preference Alignment
 Abstract.
@@
 To overcome this limitation and develop more reliable LLMs for recommendations, we propose a new Deliberative Recommendation task, which incorporates explicit reasoning about user preferences as an additional alignment goal.
-We then introduce the Deliberative User Preference Alignment framework, designed to enhance reasoning capabilities by utilizing verbalized user feedback in a step-wise manner to tackle this task.
+We then introduce the Reasoning-powered Recommender framework for deliberative user preference alignment, designed to enhance reasoning capabilities by utilizing verbalized user feedback in a step-wise manner to tackle this task.
 The framework employs collaborative step-wise experts and tailored training strategies for each expert.
@@
 In this way, we highlight three fundamental capabilities: 1) Preference Distillation, which analyzes the user history to identify aspect-level user preferences (Guan et al., 2019) and examines existing verbalized user feedback on the candidate item to recognize positive and negative features; 2) Preference Matching, which matches distilled user preferences with item features and generates rationales for why the user might like or dislike the candidate item; and 3) Feedback Prediction, which predicts the user feedback with consideration of the generated raionales.
-In this light, we propose a Deliberative User Preference Alignment (DeliRec) framework to teach LLMs with the recognized reasoning capabilities with the guidance of verbalized user feedback (e.g., user reviews (Hasan et al., 2024)).
-As illustrated in Figure 2, DeliRec employs three collaborative experts, each dedicated to a specific reasoning step, to hone step-specific reasoning capabilities.
+In this light, we propose a Reasoning-powered Recommender (Reason4Rec) framework to teach LLMs with the recognized reasoning capabilities with the guidance of verbalized user feedback (e.g., user reviews (Hasan et al., 2024)).
+As illustrated in Figure 2, Reason4Rec employs three collaborative experts, each dedicated to a specific reasoning step, to hone step-specific reasoning capabilities.
 To reduce the computational and memory cost, these experts are implemented using separate QLoRA (Dettmers et al., 2023) adapters on the same LLM.
 Recognizing that each reasoning step necessitates extracting distinct signals from the verbalized user feedback, each expert is equipped with a uniquely tailored training strategy.
-Taking rating prediction (Liu et al., 2019a) as an example, we conduct extensive experiments across three real-world datasets, validating the effectiveness of DeliRec regarding prediction accuracy and reasoning quality, and the rationality of the framework design.
-The code and datasets are available at https://anonymous.4open.science/r/OurDeliRec-F4B5.
+Taking rating prediction (Liu et al., 2019a) as an example, we conduct extensive experiments across three real-world datasets, validating the effectiveness of Reason4Rec regarding prediction accuracy and reasoning quality, and the rationality of the framework design.
+The code and datasets are available at https://github.com/Peter-Fy/Reason4Rec.
 To summarize, our contributions are threefold:
 We formulate the Deliberative Recommendation task, which pursues LLMs conducting reasoning before making prediction by learning from verbalized user feedback.
-We propose a Deliberative User Preference Alignment framework, which achieves the reasoning process with step-wise experts associated with specifically designed training strategies.
-We conduct extensive experiments on three datasets, validating the effectiveness and rationality of the proposed DeliRec framework, showing the potential of slow thinking in recommendation.
+We propose a Reasoning-powered Recommender framework for deliberative user preference alignment, which achieves the reasoning process with step-wise experts associated with specifically designed training strategies.
+We conduct extensive experiments on three datasets, validating the effectiveness and rationality of the proposed Reason4Rec framework, showing the potential of slow thinking in recommendation.
 2. TASK FORMULATION
@@
 3. METHOD
-In this section, we present DeliRec with three experts to collaboratively perform multi-step reasoning before recommendation.
-To supervise the reasoning process, DeliRec incorporates user reviews, directly aligning the reasoning with user preferences rather than only inferring preferences through users’ numerical or binary feedback (e.g., ratings and clicks).
+In this section, we present Reason4Rec with three experts to collaboratively perform multi-step reasoning before recommendation.
+To supervise the reasoning process, Reason4Rec incorporates user reviews, directly aligning the reasoning with user preferences rather than only inferring preferences through users’ numerical or binary feedback (e.g., ratings and clicks).
 We then detail the three experts, i.e., summarizer, reasoner, and predictor, with their fine-tuning strategies.
-An overview of DeliRec is illustrated in Figure 2.
+An overview of Reason4Rec is illustrated in Figure 2.
 3.1. Summarizer for Preference Distillation
-DeliRec leverages a summarizer to distill the user’s preferences and the item’s positive and negative features from their historical reviews.
+Reason4Rec leverages a summarizer to distill the user’s preferences and the item’s positive and negative features from their historical reviews.
 It filters out noise from the reviews, captures nuanced user preferences, and enriches item features with the LLM’s inherent knowledge.
@@
 3.2. Reasoner for Preference Matching
-Given all the aspect-preference summaries of a user and an item, DeliRec adopts a reasoner to evaluate their matching degree, generating reasons for why the user might like or dislike the item.
+Given all the aspect-preference summaries of a user and an item, Reason4Rec adopts a reasoner to evaluate their matching degree, generating reasons for why the user might like or dislike the item.
 Reasoner Instantiation.
@@
 3.3. Predictor for Feedback Prediction
-DeliRec finally introduces a predictor for user feedback prediction.
+Reason4Rec finally introduces a predictor for user feedback prediction.
 In this work, we investigate the classic rating prediction task (Zheng et al., 2017).
@@
 In this section, we conduct a series of experiments to answer the following research questions:
-RQ1: How does DeliRec perform compared to other baseline methods in terms of accuracy in predicting user feedback (i.e., ratings) and the quality of generated reasons?
-RQ2: Why is the multi-step reasoning strategy in DeliRec essential, and how does it compare in effectiveness to one-step reasoning strategies?
+RQ1: How does Reason4Rec perform compared to other baseline methods in terms of accuracy in predicting user feedback (i.e., ratings) and the quality of generated reasons?
+RQ2: Why is the multi-step reasoning strategy in Reason4Rec essential, and how does it compare in effectiveness to one-step reasoning strategies?
 RQ3: How effective is our approach (the reason generation method and proposed reward model) in incorporating user-verbalized preference feedback?
@@
 4.1.3. Baselines.
-To comprehensively evaluate performance, we compare DeliRec against three categories of recommendation methods: traditional methods (MF), review-based recommendation methods (DeepCoNN, NARRE, DAML), and LLM-based methods (GPT-4o, Rec-SAVER, EXP3TR).
+To comprehensively evaluate performance, we compare Reason4Rec against three categories of recommendation methods: traditional methods (MF), review-based recommendation methods (DeepCoNN, NARRE, DAML), and LLM-based methods (GPT-4o, Rec-SAVER, EXP3TR).
 MF (Koren et al., 2009): Matrix Factorization is a classical collaborative filtering method that predicts user ratings based solely on the historical rating matrix.
@@
 from previous work (Liu et al., 2019b).
-For all reasoning-enhanced LLMRec methods (i.e., Rec-SAVER, EXP3RT and our DeliRec), we use GPT-3.5-turbo666https://platform.openai.com/docs/models#gpt-3-5-turbo.
+For all reasoning-enhanced LLMRec methods (i.e., Rec-SAVER, EXP3RT and our Reason4Rec), we use GPT-3.5-turbo666https://platform.openai.com/docs/models#gpt-3-5-turbo.
 as the teacher model and fine-tune the LLama3-8B model777https://www.llama.com/docs/model-cards-and-prompt-formats/meta-llama-3/.
@@
 4.2. Main Results (RQ1)
-We compare our DeliRec method with baseline methods on the accuracy of user feedback (rating) prediction (Predictor output) and the quality of generated reasons (Reasoner output).
+We compare our Reason4Rec method with baseline methods on the accuracy of user feedback (rating) prediction (Predictor output) and the quality of generated reasons (Reasoner output).
 Accuracy of User Feedback Prediction.
@@
 For the RMSE metric, our method achieves the best performance on the Amazon Music and Amazon Book datasets, while delivering comparable results to the best baseline on the Yelp dataset.
-These superior performances of DeliRec confirm that incorporating reasoning about user preference and item features before making the final prediction, i.e., applying deliberation, is effective in enhancing RecLLM performance, thereby validating the rationality of our formulation of deliberative recommendation.
+These superior performances of Reason4Rec confirm that incorporating reasoning about user preference and item features before making the final prediction, i.e., applying deliberation, is effective in enhancing RecLLM performance, thereby validating the rationality of our formulation of deliberative recommendation.
 Notably, although Rec-SAVER and EXP3RT also incorporate reasoning, their reasoning is performed in a single step, unlike our approach, which involves multiple steps—preference distillation, matching, and prediction.
@@
 Quality of Generated Reasons.
-We next analyze the quality of the reasons generated by the Reasoner in DeliRec, compared to two reasoning-aware baselines (i.e., Rec-SAVER and EXP3RT).
+We next analyze the quality of the reasons generated by the Reasoner in Reason4Rec, compared to two reasoning-aware baselines (i.e., Rec-SAVER and EXP3RT).
 Table 3 summarizes the results.
@@
 | EXP3RT | 0.5608 | 0.8385 | 0.6135 | 0.9370 | 0.8306 | 1.2311 | |
-| Ours | DeliRec | 0.5442 | 0.7722 | 0.6029 | 0.8345 | 0.7586 | 1.0418 |
+| Ours | Reason4Rec | 0.5442 | 0.7722 | 0.6029 | 0.8345 | 0.7586 | 1.0418 |
 | Method | Music | Book | Yelp | |||
@@
 | EXP3RT | 76.22 | 0.3840 | 73.60 | 0.4373 | 64.28 | 0.4275 |
-| DeliRec | 80.53 | 0.4067 | 77.31 | 0.4731 | 72.70 | 0.4565 |
+| Reason4Rec | 80.53 | 0.4067 | 77.31 | 0.4731 | 72.70 | 0.4565 |
 4.3. In-depth Analysis
-In the following sections, we present an analysis of the impact of multi-step reasoning and verbalized preference feedback in DeliRec, followed by the evaluation of DeliRec’s inference cost.
-Lastly, we provide detailed cases comparing the reasoning generated by DeliRec with those from baseline methods to demonstrate its superior alignment with the user’s true preferences.
+In the following sections, we present an analysis of the impact of multi-step reasoning and verbalized preference feedback in Reason4Rec, followed by the evaluation of Reason4Rec’s inference cost.
+Lastly, we provide detailed cases comparing the reasoning generated by Reason4Rec with those from baseline methods to demonstrate its superior alignment with the user’s true preferences.
 4.3.1. Impact of Multi-step Reasoning (RQ2)
-To evaluate the impact of multi-step reasoning in DeliRec, we first conduct an ablation study to analyze the necessity of the component in each step (i.e., Summarizer, Reasoner and Predictor) and then assess the performance changes by replacing the multi-step reasoning mechanism with two alternative single-step reasoning strategies.
+To evaluate the impact of multi-step reasoning in Reason4Rec, we first conduct an ablation study to analyze the necessity of the component in each step (i.e., Summarizer, Reasoner and Predictor) and then assess the performance changes by replacing the multi-step reasoning mechanism with two alternative single-step reasoning strategies.
 Necessity of Each Step.
-An ablation study on DeliRec is conducted by removing each step to assess its contribution.
+An ablation study on Reason4Rec is conducted by removing each step to assess its contribution.
 Specifically, we make the following changes in each experiment: 1) w/o Step 1 (i.e., Summarizer): Directly using user historical reviews to replace the aspect-preference summaries as input for Step 2 and Step 3; 2) w/o Step 2 (i.e., Reasoner): Remove the Reason Placeholder section from the Step 3 Prompt and instruct the LLMs to directly predict the rating based solely on the aspect-preference summaries and historical ratings; 3) w/o Step 3 (i.e., Predictor): Combine Step 2 and Step 3 into a single step, asking the LLMs to simultaneously generate the match analysis and predicted ratings based on the aspect-preference summaries and the user’s historical ratings.
@@
 Multi-step.
-To verify the effectiveness of the multi-step reasoning strategy in DeliRec, we compare it with two alternative strategies: (1) One-step reasoning, which generates both an entangled reasoning process and the ratings in a single step, and (2) Chain-of-Thought (CoT) reasoning, which generates the outputs of our three reasoning steps in one single prompt.
+To verify the effectiveness of the multi-step reasoning strategy in Reason4Rec, we compare it with two alternative strategies: (1) One-step reasoning, which generates both an entangled reasoning process and the ratings in a single step, and (2) Chain-of-Thought (CoT) reasoning, which generates the outputs of our three reasoning steps in one single prompt.
 We prepare the datasets separately based on the two above strategies and train the LLMs accordingly.
@@
 This may be due to the complexity of multiple tasks involved in the CoT reasoning strategy, which poses significant challenges for LLMs to effectively process simultaneously.
+| Method | Music | Book | ||
+| MAE | RMSE | MAE | RMSE | |
+| Without Reason | 0.5529 | 0.7742 | 0.6225 | 0.8394 |
+| Review-guide Reason | 0.5825 | 0.7812 | 0.6376 | 0.8390 |
+| Review-inferred Reason | 0.5723 | 0.7816 | 0.6306 | 0.8420 |
+| History-inferred Reason | 0.5640 | 0.7744 | 0.6305 | 0.8376 |
+| Reason4Rec | 0.5442 | 0.7722 | 0.6029 | 0.8345 |
 4.3.2. Effect of Reason Generation Approaches (RQ3)
-To validate the effectiveness of our reason generation approach in Step 2 of DeliRec, we compare it against four alternative approaches: 1) Without reason: Remove Step 2 from DeliRec entirely and instruct the LLMs to directly predict the rating based solely on the aspect-preference summaries and historical ratings, as done in the ablation study; 2) Review-guided reasons: Use the target item reviews as direct supervision for the LLMs’ learning process in Step 2; 3) Review-inferred reasons: Use ChatGPT to generate reasons inferred from the target item reviews and employ these generated reasons as supervision for Step 2; 4) History-inferred reason: Use ChatGPT to generate reasons inferred from the aspect-preference summaries using the Step 2 Prompt and directly use them as supervision for Step 2 without applying filtering through our reward model.
+To validate the effectiveness of our reason generation approach in Step 2 of Reason4Rec, we compare it against four alternative approaches: 1) Without reason: Remove Step 2 from Reason4Rec entirely and instruct the LLMs to directly predict the rating based solely on the aspect-preference summaries and historical ratings, as done in the ablation study; 2) Review-guided reasons: Use the target item reviews as direct supervision for the LLMs’ learning process in Step 2; 3) Review-inferred reasons: Use ChatGPT to generate reasons inferred from the target item reviews and employ these generated reasons as supervision for Step 2; 4) History-inferred reason: Use ChatGPT to generate reasons inferred from the aspect-preference summaries using the Step 2 Prompt and directly use them as supervision for Step 2 without applying filtering through our reward model.
 In Table 4, we present the results and make the following observations: 1) All four alternative approaches perform worse than our proposed method on both Music and Book datasets, highlighting the advantages of our proposed method in collecting high-quality reasons for training LLMs.
@@
 This indicates that directly supervising with GPT-generated reasons without refinement is ineffective, underscoring the importance of using verbalized preference feedback as supervision.
-| Method | Music | Book | ||
-| MAE | RMSE | MAE | RMSE | |
-| Without Reason | 0.5529 | 0.7742 | 0.6225 | 0.8394 |
-| Review-guide Reason | 0.5825 | 0.7812 | 0.6376 | 0.8390 |
-| Review-inferred Reason | 0.5723 | 0.7816 | 0.6306 | 0.8420 |
-| History-inferred Reason | 0.5640 | 0.7744 | 0.6305 | 0.8376 |
-| DeliRec | 0.5442 | 0.7722 | 0.6029 | 0.8345 |
 4.3.3. Effectiveness of Reward Model (RQ3)
@@
 | History-inferred Reason | 0.7257 | 1.071 | 0.6842 | 0.9686 |
-4.3.4. Cost Analysis
-To evaluate the efficiency of DeliRec, we compared its inference time and generation cost with those of reasoning-aware baselines (Rec-SAVER and EXP3RT).
-As shown in Table 6, despite dividing the reasoning process into three steps, DeliRec achieves comparable average inference time and generates a similar number of tokens as the baselines.
-Specifically, it requires approximately 6 seconds and generates an average of 148 tokens per prediction.
-This efficiency is primarily attributed to two factors: 1) the preference distillation step in DeliRec is preprocessed in advance, as discussed in Section 3.1, leaving only the last two steps to be executed during inference; and 2) DeliRec only decodes a single token during the final step, as explained in Section 3.3.
-These designs effectively minimize computation costs.
 | Method | Avg. Inference Time (s) | Avg. Tokens Generated |
-| DeliRec | 5.86 | 147.78 |
+| Reason4Rec | 5.86 | 147.78 |
 | Rec-SAVER | 6.43 | 175.59 |
 | EXP3RT | 5.62 | 150.74 |
+4.3.4. Cost Analysis
+To evaluate the efficiency of Reason4Rec, we compared its inference time and generation cost with those of reasoning-aware baselines (Rec-SAVER and EXP3RT).
+As shown in Table 6, despite dividing the reasoning process into three steps, Reason4Rec achieves comparable average inference time and generates a similar number of tokens as the baselines.
+Specifically, it requires approximately 6 seconds and generates an average of 148 tokens per prediction.
+This efficiency is primarily attributed to two factors: 1) the preference distillation step in Reason4Rec is preprocessed in advance, as discussed in Section 3.1, leaving only the last two steps to be executed during inference; and 2) Reason4Rec only decodes a single token during the final step, as explained in Section 3.3.
+These designs effectively minimize computation costs.
 4.3.5. Case study
-We conduct a case study to examine whether the reasons generated by DeliRec better align with true user preferences, as reflected in the target item reviews.
-We compare the reasoning results of DeliRec with those of EXP3RT and Rec-SAVER.
+We conduct a case study to examine whether the reasons generated by Reason4Rec better align with true user preferences, as reflected in the target item reviews.
+We compare the reasoning results of Reason4Rec with those of EXP3RT and Rec-SAVER.
 Figure 5 illustrates an example.
 In this case, the target item review indicates that the user values the book for its practicality and helpfulness.
-DeliRec accurately identifies these features and aligns them with the user’s preferences, as highlighted in green.
+Reason4Rec accurately identifies these features and aligns them with the user’s preferences, as highlighted in green.
 In contrast, EXP3RT incorrectly attributes the user’s preference to the storyline and characters.
@@
 First, they compress the complex reasoning process into a single step and train it jointly, which makes it challenging for LLMs to effectively learn.
-In contrast, DeliRec decomposes the reasoning process into multiple steps, training each step independently for better performance.
+In contrast, Reason4Rec decomposes the reasoning process into multiple steps, training each step independently for better performance.
 Second, they lack fine-grained supervision in training data selection, leading models to learn the reasoning patterns misaligned with users’ real preferences.
-DeliRec addresses this limitation by using verbalized user feedback as supervision, resulting in reasons that more accurately reflect users’ preferences and enhance the multi-step reasoning abilities of RecLLMs.
+Reason4Rec addresses this limitation by using verbalized user feedback as supervision, resulting in reasons that more accurately reflect users’ preferences and enhance the multi-step reasoning abilities of RecLLMs.
 5.2. Deliberative Alignment
@@
 While these methods leverage reviews to represent user preferences implicitly with embeddings, resulting in the lack of interoperability.
-In contrast, DeliRec generates the reasoning process for each step in explicit textual form, offering clearer and more interpretable explanations of user preferences.
+In contrast, Reason4Rec generates the reasoning process for each step in explicit textual form, offering clearer and more interpretable explanations of user preferences.
 Additionally, most review-based recommendation methods primarily leverage reviews as input features for generating embeddings.
-However, DeliRec also utilizes reviews as supervisory signals on the output side, aligning the LLM’s reasoning with the user’s true preferences.
+However, Reason4Rec also utilizes reviews as supervisory signals on the output side, aligning the LLM’s reasoning with the user’s true preferences.
 6. CONCLUSION AND FUTURE WORK
 In this work, we addressed the limitations of existing RecLLMs in complex scenarios by introducing a novel Deliberative Recommendation task, which emphasizes explicit reasoning before predicting user feedback.
-To achieve Deliberative Recommendation, we proposed the DeliRec framework, which enables multi-step reasoning via three collaborative experts with three core reasoning capabilities: Preference Distillation, Preference Matching, and Feedback Prediction.
+To achieve Deliberative Recommendation, we proposed the Reason4Rec framework, which enables multi-step reasoning via three collaborative experts with three core reasoning capabilities: Preference Distillation, Preference Matching, and Feedback Prediction.
 We aligned the reasoning process with users’ true preference by using verbalized user feedback, i.e., reviews.
-Through extensive experiments on three real-world datasets, DeliRec demonstrated superior performance in both prediction accuracy and reasoning quality, highlighting the significance of slow thinking in recommendation tasks.
-Despite its promising results, DeliRec is an initial attempt on Deliberative Recommendation, leaving many future directions.
+Through extensive experiments on three real-world datasets, Reason4Rec demonstrated superior performance in both prediction accuracy and reasoning quality, highlighting the significance of slow thinking in recommendation tasks.
+Despite its promising results, Reason4Rec is an initial attempt on Deliberative Recommendation, leaving many future directions.
 First, we only adopt user reviews as verbalized user feedback.
 Exploring multi-turn and diverse forms of verbalized user feedback, e.g., conversations, may further enhance the reasoning process.
-Second, although DeliRec achieves comparable or even lower inference costs than Rec-SAVER and EXP3RT, there is still significant room to improve its efficiency.
+Second, although Reason4Rec achieves comparable or even lower inference costs than Rec-SAVER and EXP3RT, there is still significant room to improve its efficiency.
 Future work could explore designing LLM acceleration algorithms tailored to Deliberative Recommendation.
-Third, it is promising to design an interactive learning paradigm between DeliRec and users.
+Third, it is promising to design an interactive learning paradigm between Reason4Rec and users.